# An analysis of Wigner’s friend in the framework of quantum mechanics based on the principle of typicality

### v1

@@ 10 Deutsch’s thought experiment
In 1985, modifying the Wigner’s friend paradox described in Section 8, Deutsch [11] proposed a thought experiment, which can, in principle, determine whether the observation by the observer $\mathcal{F}$ can lead to the reduction of state vector of the system $\mathcal{S}$ prepared initially in the state $|+\rangle$. In this paper, we make an analysis of Deutsch’s thought experiment in the framework of quantum mechanics based on the principle of typicality. Actually, in this paper, we study a well-known variant of this thought experiment, which is a simplification of the technique used in Frauchiger and Renner [23]. We call this variant just Deutsch’s thought experiment.
In this section, we review Deutsch’s thought experiment in the terminology of the conventional quantum mechanics. In the setting of Deutsch’s thought experiment, we consider a single qubit system $\mathcal{S}$ and two observers $\mathcal{F}$ and $\mathcal{D}$, as in the original Wigner’s friend paradox described in Section 8. The observer $\mathcal{F}$ performs a measurement $\mathcal{M}^{\mathcal{F}}$ over the system $\mathcal{S}$ in the exactly same manner as in the Wigner’s friend paradox. However, in Deutsch’s thought experiment [11], the observer $\mathcal{D}$ performs a measurement over the composite system $\mathcal{S}+\mathcal{F}$, and not over the observer $\mathcal{F}$ only, unlike in the case of the original Wigner’s friend paradox. For completeness and clarity, in what follows, we describe Deutsch’s thought experiment from scratch.
@@
We define the state $|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ by
@@
After the measurement $\mathcal{M}^{\mathcal{F}}$ by $\mathcal{F}$ over the system $\mathcal{S}$, the observer $\mathcal{D}$ performs a measurement $\mathcal{M}^{\mathcal{D}}$ described by the measurement operators $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ over the composite system $\mathcal{S}+\mathcal{F}$ (according to Postulate 4), where
@@
Applying Postulate 4 to the measurement $\mathcal{M}^{\mathcal{F}}$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, either the following two possibilities (i) or (ii) occurs:
@@
$|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle.$ (43)
Thus, applying Postulate 4 to the subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the outcome ++ occurs surely in the measurement $\mathcal{M}^{\mathcal{D}}$. Thus, for each $l=+,$-, we can consider the probability $P_{\mathrm{c2}}(l)$ that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $l$, where $P_{\mathrm{c2}}(+)=1$ and $P_{\mathrm{c2}}(-)=0$. On the other hand, the notion of the probability that the measurement $\mathcal{M}^{\mathcal{F}}$ gives a certain outcome is meaningless in this Case 2.
@@ 11 Analysis of Deutsch’s thought experiment based on the principle of typicality
$U_{\mathrm{I}}(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)=\sum_{a=0,1}(P_{a}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle$
@@
$U_{\mathrm{II}}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)$ $=\sum_{a=0,1}(P_{a}|+\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle,$ (45)
$U_{\mathrm{II}}(U_{\mathrm{I}}(|-\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)$ $=\sum_{a=0,1}(P_{a}|-\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\otimes|\Phi^{\mathcal{D}}[-]\rangle.$ (46)
@@
$U_{\mathrm{II}}(U_{\mathrm{I}}(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)=\sum_{a=0,1}\sum_{b=+,-}(P_{a}Q_{b}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\otimes|\Phi^{\mathcal{D}}[b]\rangle$ (47)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. We denote the set $\{0,1\}\times\{+,-\}$ by $\Omega$, and we define a finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}$ by
@@
where $I_{\mathcal{S}}$ denotes the identity operator acting on $\mathcal{H}_{\mathrm{S}}$. Thus, the finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ forms measurement operators. Then, using (47) we see that the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ result in:
$U_{\mathrm{whole}}(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle):=(U_{\mathrm{II}}\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}))(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle)=\sum_{(k,l)\in\Omega}(M_{k,l}|\Psi\rangle)\otimes|\Phi[k,l]\rangle$ (49)
@@ 11.1 Application of the principle of typicality
The unitary operator $U_{\mathrm{whole}}$ applying to the initial state $|+\rangle\otimes|\Phi_{\mathrm{init}}\rangle$ describes the repeated once of the infinite repetition of the measurements where the succession of the measurement $\mathcal{M}^{\mathcal{F}}$ and the measurement $\mathcal{M}^{\mathcal{D}}$ is infinitely repeated. It follows from (49) that the application of $U_{\mathrm{whole}}$, consisiting of the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ in this order, can be regarded as a single measurement which is described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ and whose all possible outcomes form the set $\Omega$.
@@
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiments) in the above setting. This $\omega$ is an infinite sequence over $\Omega$ consisting of records in the two observers $\mathcal{F}$ and $\mathcal{D}$ which is being generated by the infinite repetition of the measurements described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\alpha$ and $\beta$ to denote the infinite sequences over $\{0,1\}$ and $\{+,-\}$, respectively, such that $(\alpha(n),\beta(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then it follows from (50) and Corollary 4.4 that $\beta$ consists only of ++, i.e., $\beta=++++++\dotsc\dotsc$. It follows from Postulate 5 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the two observers $\mathcal{F}$ and $\mathcal{D}$, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
up to the normalization factor, where the last equality follows from (48). Thus, we have that the observer $\mathcal{D}$ always records the outcome ++ after the measurement $\mathcal{M}^{\mathcal{D}}$ in every repetition of the experiment.
In addition, it follows from (50) and Theorem 4.1 that $\alpha$ is a Martin-Löf random infinite binary sequence. Therefore, it follows from Theorem 4.2 that for every $k\in\{0,1\}$ it holds that
@@
as expected from the point of view of the conventional quantum mechanics. Thus, in our world $\alpha$, the following (i) and (ii) hold:
@@
We can then see that the above situation of (i) and (ii) is different from both Case 1 and Case 2 in Section 10, although the statements of both Case 1 and Case 2 in Section 10 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given, in particular, in the Born rule, Postulate 4 (i).
We can use Postulate 7 to determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. We use $I_{\mathcal{D}}$ to denote the identity operator acting on $\mathcal{H}_{\mathcal{D}}$. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{D}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ given by
@@
where the second, third, and last equalities follow from (39), (38), and (40), respectively. On the other hand, it is easy to see that, according to Definition 7.3, the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Note also that the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ and $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ form PVMs. Thus, it follows from Postulate 7 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
$(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$
describes the repeated once of the infinite repetition of the measurements (experiments) where the measurement $\mathcal{M}^{\mathcal{D}}$ is infinitely repeated. Let $\Theta:=\{+,-\}$. The application of $U_{\mathrm{II}}$ itself, of course, forms a single measurement, which is described by the measurement operators $\{M^{\mathcal{D}}_{l}\}_{l\in\Theta}$ and whose all possible outcomes form the set $\Theta$.
@@
$(M^{\mathcal{D}}_{l}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}[l]\rangle$
@@
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiments) in the above setting. This $\omega$ is an infinite sequence over $\Theta$ consisting of records in the observer $\mathcal{D}$ which is being generated by the infinite repetition of the measurement $\mathcal{M}^{\mathcal{D}}$ described by the measurement operators $\{M^{\mathcal{D}}_{l}\}_{l\in\Theta}$ in the above setting. Since the Bernoulli measure $\lambda_{R}$ on $\Theta^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $R$-random.
Then it follows from (54) and Corollary 4.4 that $\omega$ consists only of ++, i.e., $\omega=++++++\dotsc\dotsc$. Thus, it follows from Postulate 5 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the ‘observer’ $\mathcal{F}$ both being only measured, and the real observer $\mathcal{D}$ measuring them, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
$(M^{\mathcal{D}}_{\omega(n)}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}[\omega(n)]\rangle$ $=(M^{\mathcal{D}}_{+}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}[+]\rangle$
@@
where the last equality follows from (41) and (53). It follows that, in our world $\omega$, the observer $\mathcal{D}$ always records the value ++ after the measurement $\mathcal{M}^{\mathcal{D}}$ in every repetition of the experiment. Thus, this result recovers in a refined manner Case 2 in Section 10 which is described in terms of the conventional quantum mechanics. Regarding the results of the measurement $\mathcal{M}^{\mathcal{D}}$ for the observer $\mathcal{D}$, we have the same results as in Section 11, where both treatments are based on the principle of typicality. The important point here is that neither the result of Section 11 nor the result of this section corresponds to Case 1 in Section 10 which is described in terms of the conventional quantum mechanics.

### v2

@@ 10 Deutsch’s thought experiment
In 1985, modifying the Wigner’s friend paradox, which is described in Section 8 of this paper, Deutsch [11] proposed a thought experiment which can, in principle, determine whether the observation by the observer $\mathcal{F}$ can lead to the reduction of state vector of the system $\mathcal{S}$ prepared initially in the state $|+\rangle$. In what follows, we make an analysis of Deutsch’s thought experiment in the framework of quantum mechanics based on the principle of typicality. Actually, in this paper, we study a well-known variant of this thought experiment, which is a simplification of the technique used in Frauchiger and Renner [23]. We call this variant just Deutsch’s thought experiment.
In this section, we review Deutsch’s thought experiment in the terminology of the conventional quantum mechanics. In the setting of Deutsch’s thought experiment, we consider a single qubit system $\mathcal{S}$ and two observers $\mathcal{F}$ and $\mathcal{D}$, as in the original Wigner’s friend paradox described in Section 8. The observer $\mathcal{F}$ performs a measurement $\mathcal{M}^{\mathcal{F}}$ over the system $\mathcal{S}$ in the exactly same manner as in the Wigner’s friend paradox. However, in Deutsch’s thought experiment [11], the observer $\mathcal{D}$ performs a measurement over the composite system $\mathcal{S}+\mathcal{F}$, and not over the observer $\mathcal{F}$ only, like in the case of the original Wigner’s friend paradox. For completeness and clarity, in what follows, we describe Deutsch’s thought experiment from scratch.
@@
We define a state $|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ by
@@
After the measurement $\mathcal{M}^{\mathcal{F}}$ performed by the observer $\mathcal{F}$ over the system $\mathcal{S}$, the observer $\mathcal{D}$ performs a measurement $\mathcal{M}^{\mathcal{D}}$ described by the measurement operators $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ over the composite system $\mathcal{S}+\mathcal{F}$ (according to Postulate 4), where
@@
Applying Postulate 4 to the measurement $\mathcal{M}^{\mathcal{F}}$, either the following two possibilities (i) or (ii) occurs immediately after the measurement $\mathcal{M}^{\mathcal{F}}$:
@@
According to (39), the state of the composite system $\mathcal{S}+\mathcal{F}$ is
$|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle$ (43)
immediately after the measurement $\mathcal{M}^{\mathcal{F}}$. Thus, applying Postulate 4 to the subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the outcome ++ occurs surely in the measurement $\mathcal{M}^{\mathcal{D}}$. Thus, for each $l=+,$-, we can consider the probability $P_{\mathrm{c2}}(l)$ that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $l$, where $P_{\mathrm{c2}}(+)=1$ and $P_{\mathrm{c2}}(-)=0$. On the other hand, the notion of the probability that the measurement $\mathcal{M}^{\mathcal{F}}$ gives a certain outcome is meaningless in this Case 2.
@@ 11 Analysis of Deutsch’s thought experiment based on the principle of typicality
$U_{\mathrm{I}}(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)=\sum_{a=0,1}(P_{a}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle$ (45)
@@
$U_{\mathrm{II}}((U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)$ $=\sum_{a=0,1}(P_{a}|+\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle,$ (46)
$U_{\mathrm{II}}((U_{\mathrm{I}}(|-\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)$ $=\sum_{a=0,1}(P_{a}|-\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\otimes|\Phi^{\mathcal{D}}[-]\rangle.$ (47)
@@
$U_{\mathrm{II}}((U_{\mathrm{I}}(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)=\sum_{a=0,1}\sum_{b=+,-}(P_{a}Q_{b}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\otimes|\Phi^{\mathcal{D}}[b]\rangle$ (48)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. Actually, using (46) and (47) we can check that (48) holds certainly in the case where $|\Psi\rangle=|+\rangle$ or $|\Psi\rangle=|$-$\rangle$. Thus, due to linearity we have that (48) holds for an arbitrary $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, as desired.
We denote the set $\{0,1\}\times\{+,-\}$ by $\Omega$, and we define a finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}$ by
@@
where $I_{\mathcal{S}}$ denotes the identity operator acting on $\mathcal{H}_{\mathrm{S}}$. Thus, the finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ satisfies the completeness equation, and therefore forms measurement operators. We denote the identity operator acting on $\mathcal{H}_{\mathrm{D}}$ by $I_{\mathcal{D}}$. Then, using (48) we see that the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ result in:
$U_{\mathrm{whole}}(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle):=(U_{\mathrm{II}}\circ(U_{\mathrm{I}}\otimes I_{\mathcal{D}}))(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle)=\sum_{(k,l)\in\Omega}(M_{k,l}|\Psi\rangle)\otimes|\Phi[k,l]\rangle$ (50)
@@ 11.1 Application of the principle of typicality
The unitary operator $U_{\mathrm{whole}}$ applying to the initial state $|+\rangle\otimes|\Phi_{\mathrm{init}}\rangle$ describes the repeated once of the infinite repetition of the measurements where the succession of the measurement $\mathcal{M}^{\mathcal{F}}$ and the measurement $\mathcal{M}^{\mathcal{D}}$ is infinitely repeated. It follows from (50) that the application of $U_{\mathrm{whole}}$, consisting of the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ in this order, can be regarded as a single measurement which is described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ and whose all possible outcomes form the set $\Omega$.
@@
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiment) in the above setting. This $\omega$ is an infinite sequence over $\Omega$ consisting of records in the two observers $\mathcal{F}$ and $\mathcal{D}$ which is being generated by the infinite repetition of the measurement described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\alpha$ and $\beta$ to denote the infinite sequences over $\{0,1\}$ and $\{+,-\}$, respectively, such that $(\alpha(n),\beta(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then, since $\omega$ is Martin-Löf $P$-random, it follows from (51) and Corollary 4.4 that $\beta$ consists only of ++, i.e., $\beta=++++++\dotsc\dotsc$. Thus, it follows from Postulate 5 and (50) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the two observers $\mathcal{F}$ and $\mathcal{D}$, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
up to the normalization factor, where the last equality follows from (49). Hence, we have that the observer $\mathcal{D}$ always records the outcome ++ after the measurement $\mathcal{M}^{\mathcal{D}}$ in every repetition of the experiment.
In addition, since $\omega$ is Martin-Löf $P$-random, it follows from Theorem 4.1 and (51) that $\alpha$ is a Martin-Löf random infinite binary sequence. Therefore, it follows from Theorem 4.2 that for every $k\in\{0,1\}$ it holds that
@@
where $N_{k}(\alpha\!\!\restriction_{n})$ denotes the number of the occurrences of $k$ in the prefix of $\alpha$ of length $n$. Thus, in our world $\alpha$, the following (i) and (ii) hold:
@@
We can then see that the above situation of (i) and (ii) is different from both Case 1 and Case 2 in Section 10, although the statements of both Case 1 and Case 2 in Section 10 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given.
We can use Postulate 7 to determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{D}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ in the repeated once of the experiments results in a state $|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ given by
@@
where the second, third, and last equalities follow from (39), (38), and (40), respectively. On the other hand, it is easy to see that, according to Definition 7.3, the final states of the observer $\mathcal{F}$ are not unchanged after the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Note also that each of the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ and $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ forms a PVM. Thus, it follows from Postulate 7 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
$(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$
describes the repeated once of the infinite repetition of the measurement (experiment) where the measurement $\mathcal{M}^{\mathcal{D}}$ is infinitely repeated. Let $\Theta:=\{+,-\}$. The application of $U_{\mathrm{II}}$ itself, of course, forms a single measurement, which is described by the measurement operators $\{M^{\mathcal{D}}_{l}\}_{l\in\Theta}$ and whose all possible outcomes form the set $\Theta$.
@@
$(M^{\mathcal{D}}_{l}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)))\otimes|\Phi^{\mathcal{D}}[l]\rangle$
@@
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurement (experiment) in the above setting. This $\omega$ is an infinite sequence over $\Theta$ consisting of records in the observer $\mathcal{D}$ which is being generated by the infinite repetition of the measurement $\mathcal{M}^{\mathcal{D}}$ described by the measurement operators $\{M^{\mathcal{D}}_{l}\}_{l\in\Theta}$ in the above setting. Since the Bernoulli measure $\lambda_{R}$ on $\Theta^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $R$-random.
Then, since $\omega$ is Martin-Löf $R$-random, it follows from (55) and Corollary 4.4 that $\omega$ consists only of ++, i.e., $\omega=++++++\dotsc\dotsc$. Thus, it follows from Postulate 5 and (44) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the ‘observer’ $\mathcal{F}$ both being only measured, and the real observer $\mathcal{D}$ measuring them, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
$(M^{\mathcal{D}}_{\omega(n)}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)))\otimes|\Phi^{\mathcal{D}}[\omega(n)]\rangle$ $=(M^{\mathcal{D}}_{+}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)))\otimes|\Phi^{\mathcal{D}}[+]\rangle$
@@
where the last equality follows from (41) and (54). It follows that, in our world $\omega$, the observer $\mathcal{D}$ always records the value ++ after the measurement $\mathcal{M}^{\mathcal{D}}$ in every repetition of the experiment. Thus, this result recovers in a refined manner Case 2 in Section 10 which is described in terms of the conventional quantum mechanics. Regarding the results of the measurement $\mathcal{M}^{\mathcal{D}}$ performed by the observer $\mathcal{D}$, we have the same results as in Section 11, where both of the treatments are based on the principle of typicality. The important point here is that neither the result of Section 11 nor the result of this section corresponds to Case 1 in Section 10, which is described in terms of the conventional quantum mechanics.